# Implemetation of a shooting technique for quantum optimal control on spin qudits

### 23 Mar 2026 v1

@@ Abstract
High-fidelity control of quantum systems is essential for scalable quantum technologies. We introduce a shooting-based quantum optimal control algorithm for systems of finite-dimensional Hilbert spaces, and demonstrate its performances through numerical simulations on systems inspired from single molecule magnets. Our method efficiently decomposes quantum gates into selective electromagnetic pulses, outperforming the standard GRAPE method against which it is benchmarked, especially in larger Hilbert spaces.
@@ I Introduction
One of the key challenges in advancing quantum simulations and computations is scaling computational power without compromising robustness against noise. Quantum optimal control provides a framework for steering quantum systems toward desired target states with high fidelity despite physical constraints and noise [10, 22]. By shaping external control fields, these methods enable precise manipulation of quantum dynamics. In this article, we provide the description and a benchmark for a quantum optimal control algorithm based on a gradient descent and a shooting technique [4, 28], on a concrete physical system: molecular spin qudits. By combining one or more transition metals or lanthanide ions with total spin greater than 1/2, it is possible to generate multiple quantum states. In appropriately designed molecules, these states are non-equidistant, allowing selective transitions to be driven with microwave pulses [7]. In this work, we focus on arrangements of Terbium ions between layers (called “deckers” in the literature) of phthalocyanine (Pc) ligand [20]. Though, this study extends to any process involving an addressable Hilbert space of finite dimension, for instance nitrogen-vacancy in diamond centers [17], NMR systems, trapped ions, Rydberg atoms, etc….
To implement quantum gates on the aforementioned systems, we have to unravel them into a sequence of “pulses”, which are electromagnetic signals sent to the system to control it. Multiple algorithms allow for doing this with varying efficiency. We will use the examples of the double and triple decker molecules, $\text{TbPc}_{2}$ and $\text{Tb}_{2}\text{Pc}_{3}$, to showcase our algorithm, which gives better results than the usual Gradient Ascent Pulse Engineering [21] (GRAPE) method, especially for at relatively high dimensions of the system Hilbert space. We choose comparing results to the GRAPE method since it can be used without assuming specialized knowledge on the system and is simply using a gradient descent optimization at its core, like the implementation of our method we present here. Other methods exist, for instance derived from Krotov work [23, 12], which is compared to GRAPE and other methods in [24].
@@ II.1 Quantum optimal control of pulses
In several platforms, quantum optimal control of pulses to shape quantum gates are made out of Arbitrary Wawe Generator (AWG), that can precisely control a frequency $\omega$ and a phase $\theta$ for a certain duration $\tau$. Using $\Theta$ as the Heaviside function, we can describe the outputed signal as,
@@
$\tau_{n,i}$ and $t_{n,i}$ being respectively the duration and the starting time of the $\text{i}^{th}$ train of pulsation $\omega_{n}$ with amplitude $A_{n,i}$ and phase $\theta_{n,i}$. For the a single train of pulsation $\omega$ between two levels of energy. We consider a simple system of two levels, with a drift Hamiltonian $H_{0}$ evolving according to a precession $\omega_{\text{prec}}$ around the z-axis. We can represent the AWG effect by adding a monochromatic electromagnetic field (in a direction orthogonal to the precession, the x-axis for instance) interacting with the magnetic moment of the system, which can be described by the following Hamiltonian:
@@
$=\frac{\omega_{\text{prec}}}{2}\sigma_{z}+\Omega_{\text{Rabi}}\cos(\omega t-\theta)\sigma_{x}$
@@
Usually three simplifications are made, choosing $\omega_{\text{rot}}=\omega_{\text{prec}}$ to remove the drift term which places us in the interaction picture, choosing $\omega=\omega_{\text{rot}}$, which places us in the rotating frame, and applying the Rotating Wave Approcimation (RWA) One ends up with this expression of the Hamiltonian:
@@
In the following of the paper, we place ourselves in this situation where we can apply eq. 4 between two energy levels of the Hilbert spaces we are working with as long as there are considered “connected”, that is as long as such an interaction is experimentally allowed. We can control both $\Omega_{\text{Rabi}}$ and $\theta$ through an AWG.
@@ II.1.1 An analytical method: the Givens Rotation Decomposition (GRD)
The Givens Rotation Decomposition [9] (GRD) is a method that systematically eliminates subdiagonal elements by applying succesive rotation matrices operating on some two-dimensional subspace to decompose any matrix into a “QR” decompostion [8] with Q unitary and R upper triangular. For a unitary matrix this amounts to decomposing the matrix into Q unitary and R diagonal. A Givens Rotation (GR), defined here,
$GRD(\phi,\theta)=e^{-i(\cos(\theta)\sigma_{x}+\sin(\theta)\sigma_{y})\phi}$ $=\begin{bmatrix}\cos(\phi)&-ie^{-i\theta}\sin(\phi)\\ -ie^{i\theta}\sin(\phi)&\cos(\phi)\end{bmatrix},\quad\forall\phi,\theta\in\mathbb{R},$ (5)
is used to zero out an element of a column (See the first step in eq. 6), and succesive rotations allow to zero out all subdiagonal elements for a given column (The second step in eq. 6). Then we can do it on the next column, without removing the zeros on the previous columns since the GR applied on zeros does not modify them.
$U\xrightarrow[]{}\left(\begin{array}[]{cccc}u_{1,1}&u_{1,2}&\cdots&u_{1,N}\\ \vdots&\vdots&\ddots&\vdots\\ u_{N-2,1}&u_{N-2,2}&\cdots&u_{N-2,N}\\ \tilde{u}_{N-1,1}&\tilde{u}_{N-1,2}&\cdots&\tilde{u}_{N-1,N}\\ 0&\tilde{u}_{N,2}&\cdots&\tilde{u}_{N,N}\end{array}\right)\xrightarrow[]{}\left(\begin{array}[]{cccc}\tilde{u}_{1,1}&\tilde{u}_{1,2}&\cdots&\tilde{u}_{1,N}\\ 0&\tilde{\tilde{u}}_{2,2}&\cdots&\tilde{\tilde{u}}_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ 0&\tilde{\tilde{u}}_{N-1,2}&\cdots&\tilde{\tilde{u}}_{N-1,N}\\ 0&\tilde{u}_{N,2}&\cdots&\tilde{u}_{N,N}\end{array}\right)\xrightarrow[]{}\left(\begin{array}[]{cccc}\tilde{u}_{1,1}&\tilde{u}_{1,2}&\cdots&\tilde{u}_{1,N}\\ 0&\tilde{\tilde{\tilde{u}}}_{2,2}&\cdots&\tilde{\tilde{\tilde{u}}}_{2,N}\\ 0&0&\ddots&\vdots\\ 0&0&\cdots&\tilde{\tilde{\tilde{u}}}_{N-1,N}\\ 0&0&\cdots&\tilde{\tilde{\tilde{u}}}_{N,N}\end{array}\right)$ (6)
@@
We can make the z rotations needed to remove the $e^{-i\xi_{i}}$ using some combinations of x and y rotations, to end up with the identity matrix at the right of eq. 7, which means the applied operations would have constructed $U^{\dagger}$ (up to a global phase). Thus by inputing the transconjugate of a target unitary matrix $U$ in the GRD, we can construct $U$. An important point of eq. 5 is that using the Hamiltonian in eq. 4 to make a GR, the amplitude $\Omega$ of a pulse $e^{-iHt}$ and its duration $t$ are equivalent and represented by the angle $\phi$. Thus a longer pulse of a smaller amplitude is equivalent to a shorter pulse of a larger amplitude as long as their product are equal. This would not be the case if we were to consider a drift term. In practice the experimentators cannot go higher than a certain amplitude, which would heat up the system too much and cause the excitation of unwanted levels according to [7]. Moreover, mathematically, a higher amplitude pulse would induce a non-negligeable Bloch-Siegert effect [3] due to the fast-oscillating terms neglected in the RWA, which would make eq. 4 not valid anymore. Thus it is not possible to make arbitrarily faster quantum gates just by considering pulses of larger amplitudes.
One important drawback of the GRD is that it uses only monochromatic pulses, targeting one transition at a time. Optimization methods give better results, and are therefore preferred in practice.
@@ II.1.2 An optimization algorithm: the GRadient Ascent Pulse Engineering (GRAPE) method
The GRAPE method [21] decomposes a quantum operation into a sequence of multi chromatic pulses and follows a gradient descent described in [21, Section 2]. For numerical results in section III.2 we will use the GRAPE method from the [18] library since this this package is widely used. Assume one can reduce a physical system to an Hamiltonian composed of available control terms $H_{k}$ with $k\leq N_{\text{ctrls}}\in\mathbb{N}$ controlled by some continuous functions $u_{k}(t)$. We want to reach an objective gate $O$ by constructing a gate $U$ as close as possible and defined by (considering $\hbar=1$):
@@
Because optimizing functions of time, the controls $u_{k}(t)$, is difficult since they live in an infinite dimensional space, we discretize the problem. $O$ has to be reached with a finite sequence of $N$ gates $U_{j}$ with $j\leq N\in\mathbb{N}$ that write as:
$U_{j}$ $=e^{-i\Delta t(H_{0}+\sum_{k=1}^{N_{\text{ctrls}}}u_{j,k}(t)H_{k})},$
@@
With dim being the dimension the Hilbert space at hand, the cost function is considered to be
$\text{cost}=1-\left|\text{Tr}\left(O^{\dagger}U\right)\right|/\text{dim}.$ (9)
We will refer to it as “the infidelity”, since it is referenced as such in [18], although other definitions using this name exist. To update the control terms $u_{j,k}(t)$ different techniques can be used. QuTiP [18] uses the L-BFGS-B [6] method of SciPy [33] for instance.
II.1.3 Comparing implementation times between methods
We chose to compare implementations times for the optimization methods shown in this paper (such as GRAPE) by bounding the controls outputted by the algorithm (and possibly much higher than physically allowed) like in the GRD. The GRD uses monochromatic pulses that have the same amplitude $\Omega$ (the maximum possibly available given experimental constraints) for different amount of time. That is we determine a $\theta$ and a $\phi$ from eq. 5, and we choose a time $t$ such that $\phi=\Omega t$. However the GRAPE method uses discretized multichromatic pulses over a fixed amount of time with different amplitudes, some possibly higher than what the AWG can output. Looking at the frequence $i$ with $\sigma_{x,i}$ and $\sigma_{y,i}$ the x and y transitions for this frequence, $\alpha_{i}$ and $\beta_{i}$ being unbounded real numbers, to compare optimal methods to the analytical GRD algorithm, we will need to translate the Hamiltonian applied over each discretized time $\Delta t$ in GRAPE such a way:
$H_{GRAPE}$ $=\sum_{i=1}^{N}\left[\alpha_{i}\sigma_{x,i}+\beta_{i}\sigma_{y,i}\right]$
$=\sum_{i=1}^{N}\left[\alpha_{i}\sigma_{x,i}+\beta_{i}\sigma_{y,i}\right]=\Delta t\sum_{i=1}^{N}\Omega_{i}\left[\cos(\theta_{i})\sigma_{x,i}+\sin(\theta_{i})\sigma_{y,i}\right].$ (10)
@@ II.1.3 Comparing implementation times between methods
$\Delta t$ $=\sum_{j=1}^{N}\Omega_{j},$ (11)
@@
$\theta_{i}$ $=\arctan\left(\frac{\beta_{i}}{\alpha_{i}}\right).$ (13)
We have to normalize the $\Omega_{i}$ for the sum of the $\Omega_{i}$ to not go higher than $\Omega$ for experimental reasons as explained at the end of section II.1.1. We take 1 to be the $\Omega_{i}$ maximum value to make the link with a similar scaling of $\phi$ in eq. 5. This framework means that it takes $\Omega^{-1}\times\pi/2$ units of time to inverse two populations, with this $\Omega$ being the maximum available amplitude sustainable by the system. This comparison using an equivalence between time and amplitude of the inputted pulses is made possible when there is no drift term.
@@
The Method for Adjoint and Gradient-based self-Iterative Construction And Refinement of Pulses (MAGICARP) is a method developed in [15] and named as such to distinguish it from conventional gradient techniques. The core of this article lies in its numerical implementation for a system inspired from the physics of SMMs, especially for the triple and double deckers systems detailed in section III.1.

### 10 Apr 2026 v2

# Implementation of a shooting technique for quantum unitary control on spin qudits

@@ Abstract
High-fidelity control of quantum systems is essential for scalable quantum technologies. We introduce a shooting-based method which yields smooth control pulses designed to implement gates on discrete quantum systems, and demonstrate its performances through numerical simulations on systems inspired from single molecule magnets. Our method efficiently decomposes quantum gates into electromagnetic pulses, and determines control pulses which are faster than GRAPE, all the more as the system’s dimension increases.
@@ I Introduction
One of the key challenges in advancing quantum simulations and computations is scaling computational power without compromising robustness against noise. Quantum optimal control provides a framework for steering quantum systems toward desired target states with high fidelity in the presence of physical constraints and noise [13, 25]. By shaping external control fields, these methods enable precise manipulation of quantum dynamics. In this work, we provide the description and a numerical benchmark for a quantum optimal control algorithm based on a gradient descent and a shooting technique [6, 31], on a concrete physical system: molecular spin qudits. By combining one or more transition metals or lanthanide ions with total spin greater than 1/2, it is possible to generate multiple quantum states. In appropriately designed molecules, the energies of these states are non-equidistant, allowing selective transitions to be addressed by microwave pulses [10]. In this work, we focus on arrangements of terbium ions between layers (called “deckers” in the literature) of phthalocyanine (Pc) ligands [23]. However, this study may be extended to any process involving an addressable Hilbert space of finite dimension, for instance nitrogen-vacancy in diamond centers [20], NMR systems, trapped ions, Rydberg atoms.
To implement quantum gates on the aforementioned systems, we have to unravel the gates into a sequence of electromagnetic signals sent to the system. Multiple algorithms have been designed to address this issue and most notably the widely used Gradient Ascent Pulse Engineering (GRAPE). This paper showcases a shooting-based method by comparing it to GRAPE on two physically relevant examples, the double and triple decker molecules, $\text{TbPc}_{2}$ and $\text{Tb}_{2}\text{Pc}_{3}$. Our numerical results demonstrate that this shooting method, achieves the desired fidelity level and provides faster execution times, up to 37% for the triple-decker molecule in the case of a quantum Fourier transform for instance. We denote by execution time the actual time it takes to run the gate on a quantum platform and not the (classical) time needed to find the pulses. We chose to numerically compare the results of these two methods, since both can be used without assuming specialized knowledge on the system and implement gradient-based optimisations at their cores. Other methods exist, for instance derived from the work of Krotov [26, 15], which has been compared to GRAPE and other methods in [27].
@@ II.1 Quantum optimal control of pulses
In several platforms, quantum optimal control of pulses to shape quantum gates are made out of Arbitrary Waweform Generator (AWG), that can accurately control a frequency $\omega$ and a phase $\theta$ for each transition $n$, and each pulse timestep $i$ having duration $\tau$, where $A$ is the amplitude of the pulse. Using $\Theta$ as the Heaviside function, the pulse function $S(t)$ is represented here by:
@@
We consider a simple system of two levels, with a drift Hamiltonian $H_{0}$ evolving according to a precession $\omega_{\text{prec}}$ around the z-axis. We can represent the AWG effect by adding a monochromatic electromagnetic field (in a direction orthogonal to the precession, the x-axis for instance) interacting with the magnetic moment of the system [14], which can be described by the following Hamiltonian:
@@
$=\frac{\omega_{\text{prec}}}{2}\sigma_{z}+\Omega_{\text{Rabi}}\cos(\omega t-\theta)\sigma_{x}.$
@@
Usually three simplifications are made, choosing $\omega_{\text{rot}}=\omega_{\text{prec}}$ to remove the drift term which places us in the interaction picture, choosing $\omega=\omega_{\text{rot}}$, which places us in the rotating frame, and applying the Rotating Wave Approcimation (RWA) to neglect fast oscillating terms, one ends up with this expression of the Hamiltonian:
@@
In the following of the paper, we place ourselves in this situation where we can apply eq. 4 between two energy levels of the Hilbert spaces we are working with as long as they are considered “connected”, that is as long as such an interaction is experimentally allowed. We can control both $\Omega_{\text{Rabi}}$ and $\theta$ through an AWG.
@@ II.1.1 An analytical method: the Givens Rotation Decomposition (GRD)
The Givens Rotation Decomposition [12] (GRD) is a method that systematically eliminates subdiagonal elements by applying successive rotation matrices operating on some two-dimensional subspace to decompose any matrix into a “QR” decompostion [11] with $Q$ unitary and $R$ upper triangular. For a unitary matrix this amounts to decomposing the matrix into $Q$ unitary and $R$ diagonal. A Givens Rotation (GR), defined by two real parameters $\theta$ and $\phi$ as
$\text{GR}(\phi,\theta)=e^{-i(\cos(\theta)\sigma_{x}+\sin(\theta)\sigma_{y})\phi}$ $=\begin{bmatrix}\cos(\phi)&-ie^{-i\theta}\sin(\phi)\\ -ie^{i\theta}\sin(\phi)&\cos(\phi)\end{bmatrix}$, (5)
is embedded in a larger space to be used to zero out an element of a column (See the first step in eq. 11), and succesive rotations allow to zero out all subdiagonal elements for a given column (The second step in eq. 11). Then we can do it on the next column, without removing the zeros on the previous columns since the GR applied on zeros does not modify them.
$U$ $\xrightarrow[]{}\left(\begin{array}[]{cccc}u_{1,1}&u_{1,2}&\cdots&u_{1,N}\\ \vdots&\vdots&\ddots&\vdots\\ u_{N-2,1}&u_{N-2,2}&\cdots&u_{N-2,N}\\ \tilde{u}_{N-1,1}&\tilde{u}_{N-1,2}&\cdots&\tilde{u}_{N-1,N}\\ 0&\tilde{u}_{N,2}&\cdots&\tilde{u}_{N,N}\end{array}\right)$ (11)
$\xrightarrow[]{}\left(\begin{array}[]{cccc}\tilde{u}_{1,1}&\tilde{u}_{1,2}&\cdots&\tilde{u}_{1,N}\\ 0&\tilde{\tilde{u}}_{2,2}&\cdots&\tilde{\tilde{u}}_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ 0&\tilde{\tilde{u}}_{N-1,2}&\cdots&\tilde{\tilde{u}}_{N-1,N}\\ 0&\tilde{u}_{N,2}&\cdots&\tilde{u}_{N,N}\end{array}\right)\xrightarrow[]{}\left(\begin{array}[]{cccc}\tilde{u}_{1,1}&\tilde{u}_{1,2}&\cdots&\tilde{u}_{1,N}\\ 0&\tilde{\tilde{\tilde{u}}}_{2,2}&\cdots&\tilde{\tilde{\tilde{u}}}_{2,N}\\ 0&0&\ddots&\vdots\\ 0&0&\cdots&\tilde{\tilde{\tilde{u}}}_{N-1,N}\\ 0&0&\cdots&\tilde{\tilde{\tilde{u}}}_{N,N}\end{array}\right)$ (22)
@@
We can make the z rotations needed to remove the $e^{-i\xi_{i}}$ using some combinations of x and y rotations, to end up with the identity matrix at the right of eq. 23, which means the applied operations would have constructed $U^{\dagger}$ (up to a global phase). Thus by inputing the conjugate transpose of a target unitary matrix $U$ in the GRD, we can construct $U$. An important point of eq. 5 is that, using the Hamiltonian in eq. 4 to make a GR, the amplitude $\Omega$ of a pulse $e^{-iHt}$ and its duration $t$ are equivalent and represented by the angle $\phi$. Thus a longer pulse of a smaller amplitude is equivalent to a shorter pulse of a larger amplitude as long as their product are equal. This would not be the case if we were to consider a drift term. In practice the power output of the experimental devices must not exceed a certain threshold, otherwise the pulses could damage the molecule or induce undesired transitions between energy levels. [10]. Moreover, mathematically, a higher amplitude pulse would induce a non-negligible Bloch-Siegert effect [5] due to the fast-oscillating terms neglected in the RWA, which would make eq. 4 not valid anymore. Thus it is not possible to make arbitrarily faster quantum gates just by considering pulses of larger amplitudes.
The GRD algorithm relies only on linear algebra, is interpretable and terminates in a finite number of steps. However these steps must be applied sequentially. On the other hand, the techniques relying on optimization generally do not terminate in a finite number of steps and could be attracted to local extrema, but they are able to handle multi-chromatic pulses and thereby achieve significantly faster implementations of gates.
@@ II.1.2 An optimization algorithm: the GRadient Ascent Pulse Engineering (GRAPE) method
The GRAPE method [24] decomposes a quantum operation into a sequence of multi-chromatic pulses and follows a gradient descent described in [24, Section 2]. For numerical results in section III.2 we will use the GRAPE method from the QuTiP library since this package is widely used [21]. Assume one can model a physical system by an Hamiltonian composed of available control terms $H_{k}$ with $k\leq N_{\text{ctrls}}\in\mathbb{N}$ controlled by some continuous functions $u_{k}(t)$. We want to reach an objective gate $O$ by constructing a gate $U$ as close as possible and defined by (considering $\hbar=1$):
@@
Since determining optimal controls in closed-form is usually out of reach, we choose a discretization scheme and numerically solve an non-linear program (direct approach). In this case, control functions are replaced by piecewise constant functions: for each piece, the duration is fixed to $\Delta t$ but the amplitudes are optimization variables. $O$ has to be reached with a finite sequence of $N$ gates $U_{j}$ with $j\leq N\in\mathbb{N}$ that write as:
$U_{j}$ $=e^{-i\Delta t(H_{0}+\sum_{k=1}^{N_{\text{ctrls}}}u_{j,k}H_{k})},$
@@
In a Hilbert space of dimension $d$, the cost function to be minimized reflects the error between the U and the target gate, up to a global phase. With $\mathbf{u}=\{u_{j,k}\}_{j=1,\dots,N}^{k=1,\dots,N_{\text{ctrls}}}$. It is given by
$\text{Cost}(\mathbf{u})=1-\frac{1}{d}\left|\operatorname{Tr}(O^{\dagger}U)\right|.$ (25)
and referred to as infidelity, following [21]. To update the control terms $u_{j,k}$ different optimization algorithms could be used. QuTiP [21] uses the L-BFGS-B [9] method of SciPy [35] for instance.
II.1.3 Comparing execution times between methods
To ensure a fair comparison of the execution times obtained with the methods considered in this work, the sum of the amplitude of each monochromatic pulses is bounded by a fixed amplitude $\Omega$. This is inspired from the analytical GRD approach, where the control consists of monochromatic pulses with fixed amplitude $\Omega$, chosen to be the maximum value allowed by the experimental setup, and variable duration. More precisely, once $\theta$ and $\phi$ are determined from eq. 5, the pulse duration $t$ is chosen such that $\phi=\Omega t$. By contrast, optimization methods such as the GRAPE method employ discretized multichromatic pulses defined over fixed time intervals, with independently varying amplitudes that may exceed the maximum value achievable by the arbitrary waveform generator. Therefore, in order to compare the optimal-control methods with the analytical GRD protocol on equal footing, it is necessary to recast the Hamiltonian applied during each discretized time step $\Delta t$ into an equivalent form satisfying the same amplitude bound. More specifically, for a frequency component indexed by $i$, with control operators $\sigma_{x,i}$ and $\sigma_{y,i}$, and with corresponding real-valued amplitudes $\alpha_{i}$ and $\beta_{i}$, the Hamiltonian during a time step $\Delta t$ is transformed as follows:
$H$ $=\sum_{i=1}^{N_{\text{ctrls}}}\left[\alpha_{i}\sigma_{x,i}+\beta_{i}\sigma_{y,i}\right]$
$=\sum_{i=1}^{N_{\text{ctrls}}}\left[\alpha_{i}\sigma_{x,i}+\beta_{i}\sigma_{y,i}\right]=\Delta t\sum_{i=1}^{N_{\text{ctrls}}}\Omega_{i}\left[\cos(\theta_{i})\sigma_{x,i}+\sin(\theta_{i})\sigma_{y,i}\right].$ (26)
@@ II.1.3 Comparing execution times between methods
$\Delta t$ $=\sum_{j=1}^{N_{\text{ctrls}}}\Omega_{j},$ (27)
@@
$\theta_{i}$ $=\text{atan2}\left(\frac{\beta_{i}}{\alpha_{i}}\right).$ (29)
We have to normalize the $\Omega_{i}$ for the sum of the $\Omega_{i}$, representing the energy of the pulses, to not go higher than $\Omega$ for experimental reasons as explained at the end of section II.1.1. We take 1 to be the $\Omega$ maximum value to make the link with a similar scaling of $\phi$ in eq. 5. This framework means that it takes $\Omega^{-1}\times\pi/2$ units of time to inverse two populations, with this $\Omega$ being the maximum available amplitude sustainable by the system. The comparison we propose is based on the equivalence between shorter durations and higher amplitudes, or vice-versa, and as such is relevant only when the drift Hamiltonian can be neglected.
@@
hThe Method for Adjoint and Gradient-based self-Iterative Construction And Refinement of Pulses (MAGICARP) is a method [18] and named as such to distinguish it from conventional gradient techniques. The core of this article lies in its numerical implementation for a system inspired from the physics of SMMs, especially for the triple and double deckers systems detailed in section III.1.